OpenStudy (anonymous):

write the equation of the line with slope = 7, and containing the point (0,0) A. y = 7x B. x = 7 C. y = 7 D. y = 1/7x

sam (.sam.):

. y = 7x

OpenStudy (hoblos):

line passing by origin y=mx y=7x

Directrix (directrix):

A. y = 7x y-0 = 7(x-0) y = 7x